# MalCVE: Malware Detection and CVE Association Using Large Language Models

### v1

@@ 3.1.4. Generating Search Queries
The fourth step involves issuing a separate prompt to generate additional search queries based on the summary.. This approach allows the system to use a more powerful language model while limiting the number of input tokens by excluding the code and incorporating only the summary generated in the previous step. Figure 4 shows the prompt that is passed into the LLM. The prompt is designed in this manner to reduce the generation of vague queries that refer to obfuscated identifiers or overly generic descriptions. The generated search queries are returned in plain text format and forwarded to the next step of the analysis flow to answer RQ2.
@@ 3.1.5. Semantic Similarity Search
After the code summary and search queries have been generated, as described in Sections 3.1.3 and 3.1.4, the queries are transformed into a dense vector representation. Each query is mapped to a single vector, which is subsequently used to perform an approximate nearest neighbor (ANN) (Milvus, [n. d.]) search for each vector against a database (i.e., the Milvus database described in Section 3.2.3). Each search retrieves up to 100 entries. This reduces the risk of overlooking a relevant CVE, and increases the chance of the re-ranking step finding a better fit CVE based on exact keywords. Additionally, the results of each search are aggregated by selecting the maximum similarity score for each unique CVE. An example of the aggregation is shown in Figure 5. This ensures that a CVE is not duplicated in the final results.
@@
3.1.6. Re-ranking the Semantic Search Result
To improve the ranking quality of the CVEs identified in the previous step and address RQ2, the system applies re-ranking using the BM25 algorithm (Amati, 2018). BM25 was selected due to its open-source nature and minimal computational requirements, such as not requiring a GPU. The similarity search captures high-level semantic similarities. BM25 measures direct lexical overlap between the analyzed code and CVE descriptions. This is particularly useful for identifying whether specific libraries are referenced in the CVE descriptions.
@@ 3.1.6. Re-ranking the Semantic Search Result
When extracting CVE information, CWE identifiers associated with the CVEs are also included, when such information is available. After re-ranking, CVEs sharing the same CWE are retrieved, expanding the candidates with vulnerabilities having the same root cause.
@@ 3.1.8. Save Results to File
The final step in the analysis pipeline involves storing the prediction results, generated summary, search queries, and metadata in a file. The metadata includes information about the models and prompts utilized during the analysis.
@@ 3.2. MalCVE implementation
Figure 7 illustrates the architecture of the proof-of-concept tool MalCVE. The system includes a development and execution environment composed of GitHub Codespace and a Hetzner Cloud virtual machine (VM), which hosts the database.
@@ 3.2.1. GitHub Codespace
GitHub Codespaces are used to develop and execute a Python script that includes the Data Collector (Section 3.2.2), responsible for initializing the Milvus Vector Database (Section 3.2.3). The script also comprises the Download Worker (Section 3.2.4), Analysis Worker (Section 3.2.5), LLM Interface (Section 3.2.6), and an Evaluation module (Section 4.1).
@@ 3.2.2. Data Collector
All available technical details from the National Vulnerability Database (NVD) were stored to enable advanced querying and filtering capabilities. A notable enhancement was the addition of Common Weakness Enumeration (CWE) identifiers. By storing the CWEs associated with each CVE, the system can identify vulnerabilities that share a common root cause. This feature is particularly useful when analyzing CVEs with high similarity scores (see Section 3.1.6). Although the Common Vulnerability Scoring System (CVSS) base score and vector string are not currently utilized in the analysis, they were included to support future filtering based on severity. In the final system, the data collector stored the following data:
@@ 3.2.3. Hetzner VM and Vector Database
A Hetzner Cloud VM hosts the Milvus vector database that stores the CVEs. Hosting the database on a separate VM isolates the database from GitHub Codespaces, where malicious files are handled. The database functions as a knowledge base by incorporating relevant CVEs into the context used for the final prompt in Step 7 (see Section 3.1.7). The system is configured to periodically fetch updates from NVD, ensuring that the database remains current. This is essential for LLM to be able to accurately identify newly published CVEs.
@@ 3.2.4. Download Worker
A list of file hashes is added to the download queue by the system. The Download Worker retrieves and downloads the corresponding files from VirusShare, adhering to the platform’s rate limits. Successfully downloaded files are then transferred to the analysis queue for further processing.
@@ 3.2.5. Analysis Worker
The Analysis Worker is responsible for decompiling, deobfuscating, summarizing, generating search queries, performing semantic search, re-ranking results, generating CVE predictions, and saving the output to a file. Multiple Analysis Workers can operate concurrently. Each worker retrieves a file from the analysis queue and performs the full sequence of tasks to complete the analysis.
@@ 4.1. Evaluation Design
To evaluate RQ1, a dataset containing malware samples is required. For RQ2, a dataset that includes both malware samples and their corresponding CVE identifiers is necessary.
MalDICT (Joyce et al., 2023) was selected as the benchmarking dataset. As noted by the authors, “MalDICT includes the first public malware datasets labeled according to platform, vulnerability (CVEs), and packer” (Joyce et al., 2023). The vulnerability subset of MalDICT contains 173,886 malware samples and annotations for 128 distinct vulnerability types (Mal, [n. d.]). From this dataset, we repeatedly selected decompilable files at random until the available tokens were insufficient for further analysis. Consequently, a total of 3,839 decompilable files were selected for analysis.
Our approach is not dependent on any specific LLM. To demonstrate its effectiveness, we conducted a comprehensive benchmark of several state-of-the-art LLMs. Due to practical constraints, testing was primarily limited to models from OpenAI and Gemini. Models provided by DeepSeek were excluded, as they failed to produce structured output in accordance with the required JSON schemas, which hindered result extraction and interpretation. Additional experimentation was limited by API restrictions, latency variability, and throughput constraints, particularly with OpenAI models restricted to a rate of 200,000 tokens per minute.
For RQ1, we measured classification accuracy, i.e., what percentage of JARs were accurately classified as either malicious or benign. For RQ2, we used recall@k where (k $\in$ 1, 3, 5, 10) and accuracy as evaluation metrics. Recall@k assesses the effectiveness of the semantic similarity search by determining whether a correct CVE appears among the top-k retrieved results. Accuracy, on the other hand, evaluates the ability of the LLMs to select the correct CVE. These metrics are also commonly reported in related work (Cao and Jun, 2024), (Du et al., 2024).
@@ 4.2.1. RQ1: Malware Detection using LLMs
The dashed red lines in Figure 10 show that the mean accuracy is 97%, across all 3,839 analyzed JARs and LLMs. This corresponds to 115 misclassified files out of the total. The observed accuracy ranged from a minimum of 88% to a maximum of 100%.
The findings demonstrate that LLMs can correctly classify decompiled binaries as either malicious or benign using appropriate prompting techniques. This was achieved without any fine-tuning, relying solely on zero-shot prompting. Changes in LLM architecture did not have a significant impact on accuracy, highlighting that LLMs perform effectively with default configurations for this task.
@@ 4.2.2. RQ2: CVE association
To answer RQ2, we analyzed the performance of various LLMs to identify the best-performing systems. The number of files analyzed per model varied due to budgetary constraints, which prevented a full evaluation of all models across the entire dataset. In total, 620 JAR files were tested, distributed among the models as follows: O4-Mini (195 files), GPT-4o Mini (157 files), Gemini 2.5 Flash (125 files), Gemini 2.5 Flash Lite (86 files), and Gemini 2.5 Flash Pro (57 files). Table 1 presents classification accuracy and recall@k metrics, including their mean, maximum, minimum, and standard deviation values.
@@
To assess the impact of technical information in CVE descriptions, we included additional CVEs lacking sufficient detail in the analysis to evaluate performance differences. The results show that Recall@10 decreased by approximately 8%, Recall@5 by 11%, and Recall@3 by 6%, while Recall@1 remained largely unchanged. Accuracy also declined by 8%. These findings underscore the importance of detailed CVE descriptions for extracting relevant vulnerabilities, as LLMs are unable to reliably associate malware with CVEs when descriptions lack technical detail beyond the affected Java version. This highlights the value of publicizing and sharing comprehensive CVE information. However, expecting every CVE to include detailed technical descriptions may be impractical. Therefore, incorporating data from additional sources, such as CWEs, or crawling advisory references published by the NVD may help improve the accuracy to associate the malware with CVEs.
Impact of deobfuscation: Existing studies (Walton et al., 2025), (Xiang Li, 2023) do not account for deobfuscation. As part of a supplementary analysis, we removed the deobfuscation step in MalCVE to evaluate its impact on performance. Following this removal, Recall@10 decreased by 13%, Recall@5 by 10%, Recall@3 by 9%, and Recall@1 by 6%. Accuracy also declined by 8%. This decline is attributed to the LLM’s reduced ability to extract specific libraries and methods when deobfuscation is not applied. The search queries used in semantic similarity search are critical for retrieving relevant CVEs. Prior to deobfuscation, the extracted queries were relatively generic and frequently repeated. After applying deobfuscation, the queries became more diverse and detailed, often including specific libraries, classes, and methods (e.g., sun.awt.SunToolkit.getField, java.beans.Statement.setSecurityManager).
Summary for RQ2: MalCVE is capable of identifying CVEs assoicated with malware with a moderate level of accuracy. The absence of sufficient technical detail in CVE descriptions presents a significant barrier to achieving higher accuracy.
@@ 5. Discussion
In this section, we discuss our approach and evaluation results by comparing them with the related work, illustrating their implications for academia and industry. We also examine potential threats to the validity of our findings and describe the measures taken to address them.
@@ 5.1. Comparison with related work
Previous LLM-based malware detection studies were conducted using fewer than 250 samples. For example, Walton et al. (2025) reported results based on 200 Android APKs, while the PE ransomware study by Li, Zhu, and Zhang (Xiang Li, 2023) used an imbalanced dataset of 221 executables. In contrast, our evaluation includes 3,839 JAR files, providing greater statistical reliability. Additionally, we evaluated multiple LLMs, demonstrating the generalizability of our approach.
@@
Walton et al. (2025) explicitly excluded obfuscated APKs and, as a result, did not report data on them. In contrast, our system decompiled and applied light deobfuscation to JAR files prior to classification, achieving a mean single-class accuracy of 97% and a minimum observed accuracy of 80%, as shown in Figure 10. This level of performance is comparable to that of commercially available malware detection tools [^9] https://www.av-comparatives.org/comparison/.
@@ 5.2. Implications for Industry and Academia
Achieving 97% accuracy without fine-tuning demonstrates that generic LLMs are already viable as lightweight, scalable malware filters in real-world security workflows. At a cost of \$0.03 per file, MalCVE is significantly more cost-effective than commercial alternatives such as CrowdStrike Falcon and ANY.RUN. In the MalCVE implementation, all components, except the commercial LLMs, are open source, enabling users to replicate or customize the solution without incurring licensing fees. Security vendors can integrate the solution into existing products with minimal overhead. In particular, vendors offering analysis tools can leverage LLM-generated outputs to provide explainable verdicts to customers who require insight into why a file was flagged as malicious. Notably, this also enables small organizations to deploy a malware filter without the upfront costs associated with commercial tools.
However, a significant limitation arises from the context window constraints of LLMs. OpenAI models, including GPT-4o-mini, are unable to process large JAR binaries due to limited token capacity. Even models with extended context windows, such as Gemini 1.5 and 2.0, which support up to 1,000,000 tokens, occasionally failed to perform comprehensive semantic analysis of large binaries. As a result, complex files may be inadequately summarized or misclassified due to truncated input, thereby reducing the construct validity of the semantic representation approach. It is worth noting that newer models may not exhibit the same limitations observed in this study.
Our findings for RQ2 indicate that all evaluated models show potential for CVE association. From an industry perspective, this suggests that LLMs can support rapid analysis of suspicious files, enabling the identification of relevant CVEs and informing potential mitigation strategies. Furthermore, organizations can select models based on performance and cost considerations, making the solution accessible even to smaller organizations with limited resources. Although the current implementation focuses on Java, the approach is generalizable to other programming languages by substituting the decompiler and deobfuscator with language-specific tools.
@@ 5.3.3. Construct Validity
Filtering MalDICT to retain only vulnerabilities with detailed technical descriptions improved accuracy by reducing semantic ambiguity, but also reduced the diversity of the dataset. However, mapping malware to CVEs may not be feasible in the absence of technical information.
Malware might exploit multiple CVEs. The MalDICT dataset and our evaluation employed a single-label classification, marking correct but secondary CVEs as false negatives. This methodological limitation artificially reduces precision and may lead to an underestimation of MalCVE’s performance. However, no existing dataset offers more precise associations between malware and CVEs.
A common issue in evaluating LLM-based applications is data leakage. Since access to the MalDICT dataset requires authentication and the data consist of malware in binary format, it is unlikely that such data was used in training commercial LLMs. Therefore, we believe the risk of data leakage in our evaluation is minimal.

### v2

@@ 3.1.4. Generating Search Queries
The fourth step involves issuing a separate prompt to generate additional search queries based on the summary. This approach allows the system to use a more powerful LLM while limiting the number of input tokens by excluding the code and incorporating only the summary generated in the previous step. Figure 4 shows the prompt that is passed into the LLM. The prompt is designed to reduce the generation of vague queries that refer to obfuscated identifiers or overly generic descriptions. The generated search queries are returned in plain text format and forwarded to the next step of the analysis flow to answer RQ2.
@@ 3.1.5. Semantic Similarity Search
After the code summary and search queries have been generated, as described in Sections 3.1.3 and 3.1.4, the queries are transformed into a dense vector representation. Each query is mapped to a single vector, which is subsequently used to perform an approximate nearest neighbor (ANN) (Milvus, ) search for each vector against a database (i.e., the Milvus database described in Section 3.2.3). Each search retrieves up to 100 entries. This reduces the risk of overlooking a relevant CVE and increases the chance that the ranking step finds a better-fitting CVE based on exact keywords. Additionally, the results of each search are aggregated by selecting the maximum similarity score for each unique CVE. An example of the aggregation is shown in Figure 5. This ensures that a CVE is not duplicated in the final results.
@@
3.1.6. Ranking the Semantic Search Result
In this step, MalCVE ranks the CVEs identified in the previous step using the BM25 algorithm (Amati, 2018). BM25 was selected for its open-source nature and minimal computational requirements, including no GPU requirement. The similarity search in Step 5 captures high-level semantic similarities. BM25 measures direct lexical overlap between the analyzed code and CVE descriptions. This is particularly useful for identifying whether specific libraries are referenced in the CVE descriptions.
@@ 3.1.6. Ranking the Semantic Search Result
When extracting CVE information, CWE identifiers associated with the CVEs are included when available. After ranking, CVEs sharing the same CWE are retrieved, expanding the candidate set to include vulnerabilities with the same root cause.
@@ 3.1.8. Save Results to File
The final step in the analysis pipeline involves storing the prediction results, the generated summary, the search queries, and metadata in a single file. The metadata includes information about the models and prompts utilized during the analysis.
@@ 3.2. MalCVE implementation
Figure 7 illustrates the architecture of the proof-of-concept tool MalCVE. The system includes a development and execution environment composed of GitHub Codespace and a Hetzner Cloud virtual machine (VM) that hosts the database.
@@ 3.2.1. GitHub Codespace
GitHub Codespaces are used to develop and execute a Python script that includes the Data Collector (Section 3.2.2), which initializes the Milvus Vector Database (Section 3.2.3). The script also comprises the Download Worker (Section 3.2.4), Analysis Worker (Section 3.2.5), LLM Interface (Section 3.2.6), and an Evaluation module (Section 4.1).
@@ 3.2.2. Data Collector
All available technical details from the National Vulnerability Database (NVD) were stored to enable advanced querying and filtering capabilities. A notable enhancement was the addition of Common Weakness Enumeration (CWE) identifiers. By storing the CWEs associated with each CVE, the system can identify vulnerabilities that share a common root cause. This feature is particularly useful when analyzing CVEs with high similarity scores (see Section 3.1.6). Although the Common Vulnerability Scoring System (CVSS) base score and vector string are not currently used in the analysis, they were included to support future filtering by severity. In the final system, the data collector stored the following data:
@@ 3.2.3. Hetzner VM and Vector Database
A Hetzner Cloud VM hosts the Milvus vector database, which stores CVEs. Hosting the database on a separate VM isolates the database from GitHub Codespaces, where malicious files are handled. The database functions as a knowledge base by incorporating relevant CVEs into the context used for the final prompt in Step 7 (see Section 3.1.7). The system is configured to periodically fetch updates from NVD, keeping the database current. This is essential for an LLM to accurately identify newly published CVEs.
@@ 3.2.4. Download Worker
The system adds a list of file hashes to the download queue. The Download Worker retrieves and downloads the corresponding files from VirusShare, adhering to the platform’s rate limits. Successfully downloaded files are then transferred to the analysis queue for further processing.
@@ 3.2.5. Analysis Worker
The Analysis Worker is responsible for decompiling, deobfuscating, summarizing, generating search queries, performing semantic search, ranking results, generating CVE predictions, and saving the output to a file. Multiple Analysis Workers can operate concurrently. Each worker retrieves a file from the analysis queue and performs the full sequence of tasks to complete the analysis.
@@ 4.1. Evaluation Design
To evaluate RQ1, a dataset containing malware samples is required. For RQ2, a dataset containing both malware samples and their corresponding CVE identifiers is required.
MalDICT (Joyce et al., 2023) was selected as the benchmarking dataset. As noted by the authors, “MalDICT includes the first public malware datasets labeled according to platform, vulnerability (CVEs), and packer” (Joyce et al., 2023). The vulnerability subset of MalDICT contains 173,886 malware samples and annotations for 128 distinct vulnerability types (9). From this dataset, we repeatedly selected decompilable files at random until the available tokens within our budget were insufficient for further analysis. Consequently, a total of 3,839 decompilable files were selected for analysis.
Our approach is not dependent on any specific LLM. To demonstrate its effectiveness, we conducted a comprehensive benchmark of several state-of-the-art LLMs. Due to practical constraints, testing was primarily limited to models from OpenAI and Gemini. Models provided by De because they failed to produce structured output in accordance with the required JSON schemas, hindering result extraction and interpretation. Additional experimentation was limited by API restrictions, latency variability, and throughput constraints, particularly with OpenAI models, which were restricted to a rate of 200,000 tokens per minute.
For RQ1, we measured classification accuracy, i.e., the percentage of JARs that were correctly classified as either malicious or benign. For RQ2, we used recall@k where (k $\in$ 1, 3, 5, 10) and accuracy as evaluation metrics. Recall@k assesses the effectiveness of the semantic similarity search by determining whether a correct CVE appears among the top-k MalCVE retrieved results. Accuracy, on the otmeasures the LLMs’ ability to select the correct CVE. These metrics are also commonly reported in related work (Cao and Jun, 2024), (Du et al., 2024).
@@ 4.2.1. RQ1: Malware Detection using LLMs
The dashed red lines in Figure 10 show that the mean accuracy is 97%, across all 3,839 analyzed JARs and LLMs. This corresponds to 115 misclassified files out of the total. The observed accuracy of different LLM models ranged from 88% to 100%.
The findings demonstrate that LLMs can correctly classify decompiled binaries as either malicious or benign uswith appropriate prompting. This was achieved without any fine-tuning, relying solely on zero-shot prompting. Changes in LLM architecture did not have a significant impact on accuracy, highlighting that LLMs erform effectively with default configurations for this task.
@@ 4.2.2. RQ2: CVE association
To answer RQ2, we analyzed the performance of various LLMs to identify the best-performing systems. The number of files analyzed per model varied due to budgetary constraints, limiting a full evaluation of all models across the entire dataset. In total, 620 JAR files were tested, distributed among the models as follows: O4-Mini (195 files), GPT-4o Mini (157 files), Gemini 2.5 Flash (125 files), Gemini 2.5 Flash Lite (86 files), and Gemini 2.5 Flash Pro (57 files). Table 1 presents classification accuracy and recall@k metrics, including their mean, maximum, minimum, and standard deviation values.
@@
To assess the impact of technical information in CVE descriptions, we included additional CVEs lacking sufficient detail in the analysis to evaluate performance differences. The results show that Recall@10 decreased by approximately 8%, Recall@5 by 11%, and Recall@3 by 6%, while Recall@1 remained largely unchanged. Accuracy also declined by 8%. These findings underscore the importance of detailed CVE descriptions for extracting relevant vulnerabilities, as LLMs are unable to reliably associate malware with CVEs when descriptions lack technical detail beyond the affected Java version. This highlights the value of publicizing and sharing comprehensive CVE information. However, expecting every CVE to include detailed technical descriptions may be impractical. Therefore, incorporating data from additional sources, such as CWEs, or crawling advisory references published by the NVD, may help improve the accuracy of associating the malware with CVEs.
Impact of deobfuscation: Existing studies (Walton et al., 2025), (Xiang Li, 2023) do not account for deobfuscation. As part of a supplementary analysis, we removed the deobfuscation step in MalCVE to evaluate its impact on performance. Following this removal, Recall@10 decreased by 13%, Recall@5 by 10%, Recall@3 by 9%, and Recall@1 by 6%. Accuracy also declined by 8%. This decline is attributed to the LLM’s reduced ability to extract specific libraries and methods when deobfuscation is not applied. For example, the deobfuscator recovered critical strings indicating a call to ”findClass” that were hidden in the decompiled code. This is important because CVE-2012-4681 is defined in terms of the misuse of ”ClassFinder.findClass” and related methods. If those identifiers remain obfuscated, the model can’t connect the code to the vulnerability. The search queries used in semantic similarity search are critical for retrieving relevant CVEs. Prior to deobfuscation, the extracted queries were relatively generic and frequently repeated. After applying deobfuscation, the queries became more diverse and detailed, often including specific libraries, classes, and methods (e.g., sun.awt.SunToolkit.getField, java.beans.Statement.setSecurityManager).
Summary for RQ2: MalCVE can identify malware-associated CVEs with moderate of accuracy. The absence of sufficient technical detail in CVE descriptions presents a significant barrier to achieving higher accuracy.
@@ 5. Discussion
In this section, we discuss our approach and evaluation results, compare them with related work, and illustrate their implications for academia and industry. We also examine potential threats to the validity of our findings and describe the measures taken to address them.
@@ 5.1. Comparison with related work
Previous LLM-based malware detection studies generally operated on fewer than 250 samples. Walton et al. (Walton et al., 2025) reported results on only 200 Android APKs. The PE ransomware workwork of (Xiang Li, 2023) used an imbalanced set of 221 executables. Our evaluation covers 3,839 JAR files, yielding a much stronger statistical confidence. Additionally, we compared multiple LLM models and showed the generalizability of our approach.
@@
Walton et al. (2025) explicitly excluded obfuscated APKs and, as a result, did not report data on them. In contrast, our system decompiled and applied light deobfuscation to JAR files prior to classification, achieving a mean single-class accuracy of 97% and a minimum observed accuracy of 88%, as shown in Figure 10. This level of performance is comparable to that of commercially available malware detection tools  [^11] https://www.av-comparatives.org/comparison/.
@@ 5.2. Implications for Industry and Academia
Achieving 97% accuracy without fine-tuning demonstrates that generic LLMs are already viable as lightweight, scalable malware filters in real-world security workflows. At a cost of \$0.03 per file, MalCVE is significantly more cost-effective than commercial alternatives, such as CrowdStrike Falcon and ANY.RUN. In the MalCVE implementation, all components, except the commercial LLMs, are open source, enabling users to replicate or customize the solution without incurring licensing fees. Security vendors can integrate the solution into existing products with minimal overhead. In particular, vendors offering analysis tools can leverage LLM-generated summaries to provide explainable verdicts to customers who require insight into why a file was flagged as malicious. Notably, this also enables small organizations to deploy a malware filter without the upfront costs associated with commercial tools.
However, a significant limitation arises from the LLM context window constraints. OpenAI models, including GPT-4o-mini, cannot accommodate large JAR binaries due to limited token capacity. Even models with extensive context windows, such as Gemini 1.5 and 2.0 (1,000,000 tokens), occasionally struggled with comprehensive semantic analysis of large benign binaries. Consequently, complex files may be inadequately summarized or misclassified due to truncated input, reducing the construct validity of the semantic representation approach, although newer models might not suffer from the same issues this study encountered.
Our findings for RQ2 indicate that all evaluated models show potential for CVE association. This means that LLMs can support rapid analysis of suspicious files, enabling the identification of relevant CVEs and informing potential mitigation strategies. Furthermore, organizations can select models based on performance and cost considerations, making the solution accessible even to smaller organizations with limited resources. Although the current implementation focuses on Java, the approach is generalizable to other programming languages by substituting the decompiler and deobfuscator with language-specific tools.
@@ 5.3.3. Construct Validity
Filtering MalDICT to retain only vulnerabilities with detailed technical descriptions improved accuracy by reducing semantic ambiguity, but also reduced the dataset’s diversity. However, mapping malware to CVEs may not be feasible without technical information.
Malware might exploit multiple CVEs. The MalDICT dataset and our evaluation employed a singused single-label classification, labeling correct but secondary CVEs as false negatives. This methodological limitation artificially reduces precision and may lead to an underestimation of MalCVE’s performance. However, no other available datasets provide more precise associations between malware and CVEs.
A common issue in evaluating LLM-based applications is data leakage. Since access to the MalDICT dataset requires authentication and the data consist of malware in binary format, it is unlikely this dataset was used to train commercial LLMs. Therefore, we believe the risk of data leakage in our evaluation is minimal.